false statement by finding a counterexample

  1. x/x=1
  2. x+3)/3=x+1
  3. √(x^2+16)=x+4
1
Expert's answer
2022-05-24T14:06:13-0400

1.

If "x=0," then the quotient "\dfrac{x}{x}" is undefined.


2.

Let "x=3"


"\dfrac{3+3}{3}=2\not=4=3+1"

3.

Let "x=3"


"\sqrt{(3)^2+16}=5\not=7=3+4"